OpenStudy (anonymous):

Steps for finding the inverse of f(x)=4x-7?

OpenStudy (bibby):

the same as every other inverse problem. replace f(x) with y. exchange the y and x variables. solve for y

OpenStudy (anonymous):

I understand that, but I'm confused what to do after I get x+7 over 4.

OpenStudy (bibby):

y=4x-7 x=4y-7 x+7=4y \(\huge \frac{x+7}{4}=y\) you got it

OpenStudy (anonymous):

Oh alright, thanks :)